# Astrometric constraints on stochastic gravitational wave background with neural networks

### v1

@@ IV Results
In Fig. 2, we present the test performance results for both FCN (left panel) and GNN (right panel). The networks can perform quite well for higher values of $\Omega_{\text{GW}}$, while the predictions start to become less precise for lower values, when the noise starts to dominate. The vertical line is the theoretically estimated sensitivity for $\Omega_{\text{GW}}$ given by Eq. (3) included for reference. Note that this theoretical estimate assumes the same proper motion errors for all sources, whereas our mock data does not satisfy this assumption, as the proper motion errors depend on the magnitude of the sources. The results do not necessarily need to completely align with the theoretical line. However, we observe a tendency for the estimated $\Omega_{\text{GW}}$ values to scatter when the true value is smaller than the theoretical line, where the data is expected to be dominated by noise. As one can notice, by increasing the number of sources, the lines moves towards left, due to the inverse proportionality to the number of objects as seen in Eq. (3) and the NN predictions also improve with an increasing number of sources.
@@
Note that the plot is shown in log-scale, while our mock catalog dataset is generated with a linear-uniform distribution of $\Omega_{\text{GW}}$. We have tested both linear-uniform and log-uniform distribution of $\Omega_{\text{GW}}$ and found that NNs perform better with the linear-uniform distribution. Additionally, for the GNN, we evaluated the impact of varying distance threshold values used to define the graph connectivity for a fixed number of sources. We observed that variations in the distance threshold did not lead to substantial differences in the trained performance of the GNN.
The FCN exhibits a tendency to plateau in its predictive accuracy for smaller target values of $\Omega_{\text{GW}}$. This behavior suggests that the network struggles to capture distinctions in the lower value range, potentially due to its inherent limitations in representing complex relationships for this specific problem and the similarity of proper motion values and uncertainties to the level of noise in the data. The plateau indicates that the FCN saturates in its capacity to learn from the data with small $\Omega_{\text{GW}}$ values.
In contrast, the GNN demonstrates the ability to closely track the diagonal representing the ideal one-to-one correspondence between predicted and true values, across small values of $\Omega_{\text{GW}}$. These predictions exhibit a higher degree of dispersion for smaller $\Omega_{\text{GW}}$, which seems to be less accentuated for higher number of sources. This suggests that the GNN captures the overall trends effectively. Its dispersion appropriately increases for smaller $\Omega_{\text{GW}}$ values, reflecting its ability to adapt to the inherent uncertainties in the data.
@@
These differences can be linked to the intrinsic design of each architecture. Interestingly, for GNN, modifications of the architecture play a role between model complexity and predictive behavior. In fact, when the depth of the GNN is reduced, for instance by decreasing the number of convolutional layers to two or three, the model begins to exhibit a behavior that parallels the plateau effect seen in the FCN, as shown in Fig. 3. Despite this plateau-like behavior, the characteristic scatter of GNN predictions persists, even with fewer layers. This indicates that the scatter is not only a function of model depth but might also come from intrinsic factors such as how the GNN aggregates and propagates information within the graph structure. Although deeper GNNs are better for capturing complex dependencies, they also introduce challenges such as greater variability in predictions. These observations underscore the complex relationship between model design choices and performance outcomes, resulting in careful tuning of model architectures to balance precision and stability for specific tasks.
@@
Figure 5: The variance of the residual distributions in Fig. 4 is plotted as a function of the number of sources. In the left panel, solid lines represent the results for predictions in Fig. 4 for the FCN (yellow) and GNN (brown). Other curves correspond to results obtained by varying the number of mock data. These predictions are generated using test datasets, which constitute $20\%$ of the original mock datasets. The original mock datasets consist of $5000$ (dotted lines), $8000$ (solid lines), and $16000$ (dashdot lines) samples, respectively. The blue dashed line represents the theoretical sensitivity estimate, where the statistical error decreases proportional to $1/\sqrt{N_{s}}$, with $N_{s}$ being the number of sources. In the right panel, the variance of the residual distributions is shown by varying the range of $\Omega_{\text{GW}}$ sampling. The default range of $[0,1]$ is reduced to $[0,0.3]$ and further to $[0,0.1]$. The number of mock datasets is fixed at $8000$. Results are displayed only for the GNN.
@@
Figure 6: Predicted values of $\Omega_{\text{GW}}$ for the masked dataset are tested by changing the area of the mask. The number of sources is fixed to $N_{s}=1000$. The left panel represents the result of FCN, while the right panel the one from the GNN.
@@
Figure 7: The probability distributions of deviations between the true and predicted values for $\Omega_{\text{GW}}$ is plotted using the results obtained from the test dataset in Fig. 6, in the case of masked datasets. The cyan curve represents the Gaussian fit to the FCN results, while the blue curve represents the Gaussian fit to the GNN results.
@@
Distributions of the variance and its dependence on the source number
To further analyze the performance of the models, we plot in Fig. 4 the distribution of the residuals, i.e., the difference between the predicted and true values. We observe that the variance of the FCN is generally smaller than that of the GNN. However, the residual distribution of the FCN is more non-Gaussian and exhibits larger tails, corresponding to small $\Omega_{\text{GW}}$ realizations where the FCN displays plateau behavior in Fig. 2. For reference, we also include the Gaussian fit of the histogram in the plot.
Furthermore, we observe a clear trend where the width of the distribution decreases as the number of sources increases. To investigate this trend more explicitly, Fig.5 shows the variance as a function of the number of sources $N_{s}$. The non-smooth behavior of the curves is due to statistical fluctuations. We tested that by performing at least three runs for each point, the average profile of the trend becomes smoother, but the overall trend remains the same. For reference, we also plot a line proportional to $N_{s}^{-1/2}$, which corresponds to the theoretical prediction for astrometric sensitivity, as indicated in Eq.(3).
The left panel compares the FCN and GNN results. Consistent with Fig. 4, the FCN exhibits a smaller variance compared to the GNN. However, the variance does not decrease as $1/\sqrt{N_{s}}$ for larger numbers of sources, potentially highlighting limitations of the architecture. For the GNN, this sensitivity saturation is less pronounced but still deviates from the $1/\sqrt{N_{s}}$ scaling for large $N_{s}$. To investigate this further, we conduct additional tests by varying the number of mock data $N$, plotted in the same figure, based on the hypothesis that the requirement for more catalog realizations becomes increasingly critical as $N_{s}$ grows. We observe that the result is indeed affected when comparing the lines for $N=5000$ and $N=8000$. However, we do not find any significant difference between $N=8000$ and $N=16000$, indicating that the original number of mocks used ($N=8000$) is sufficient for our configurations.
In the right panel, we test another hypothesis that the range of $\Omega_{\text{GW}}$ for the mock data should be optimized based on the sensitivity. The theoretical sensitivity is $\Omega_{\text{GW}}\sim 0.2$ for $N_{s}=500$, while $\Omega_{\text{GW}}\sim 0.08$ for $N_{s}=12000$, but both are trained using the same range of $\Omega_{\text{GW}}=[0,1]$. As one may notice in Fig. 2, the NNs do not have many samples for noise-dominated data in the case of $N_{s}=12000$, which may explain the reduction in sensitivity. To investigate this, we test two different ranges for $\Omega_{\text{GW}}$: $\Omega_{\text{GW}}=[0,0.3]$ and $\Omega_{\text{GW}}=[0,0.1]$. We observe an improvement in sensitivity, and the $N_{s}=12000$ data point gets closer to the $1/\sqrt{N_{s}}$ curve. However, we also notice that the variance for $N_{s}=500$ decreases, likely because the samples are now taken only from small $\Omega_{\text{GW}}$ cases. While it is difficult to draw a firm conclusion from this test, it highlights the importance of the training range, as it is analogous to setting prior knowledge for the parameter values to be predicted.
Inhomogeneous case: masked samples
In the above, we have demonstrated the application of NNs with simplified mock data settings. However, there are many complex factors to consider when handling real data, and this presents a much longer path to fully realizing the potential of NNs. We leave it for the future work, while here, to illustrate one example of the flexibility of NNs, we apply them to masked data and evaluate their performance.
We fix the number of sources to $N_{s}=1000$ and mask the regions of the data set corresponding to the declination of $\delta=10^{\circ}$, $20^{\circ}$, $40^{\circ}$, and $60^{\circ}$. While we are aware that $\delta=60^{\circ}$ is non-realistic value for the mask, it is used here solely for training performance purposes. The same NN architectures are trained on these masked subsets and the results are shown in Figs. 6 and 7.
Despite the variation in input data coverage, the general trends in the results remain consistent: the FCN exhibits a plateau in performance for smaller target values, while the GNN maintains its characteristic diagonal alignment with the true values, as seen in Fig.6. The NNs successfully predict the values of $\Omega_{\rm GW}$ when the signal is dominant, but an increased dispersion is now present in the predictions of both networks, which is more clearly observed in the distributions reported in Fig.7. As expected, we observe that the variance increases with a larger mask area. This results suggests that NNs are capable of performing reasonably well even when the source distribution is inhomogeneous.
Computation time
Given the expected number of sources available in future data, a reduction in computation time is primarily anticipated for the NNs. To address this, the computational time (including the training time) is plotted in Fig. 8. All these results were evaluated on a GPU cluster. As observed, in both cases, the computation time increases logarithmically with the number of sources. The GNN exhibits higher steep in terms of time, which highlights the higher computational complexity inherent to the network itself and the manage of more complex data in input. In both cases, the behavior reflects a more accentuate computational cost for larger datasets.
Both networks are competitive in predicting the value of $\Omega_{\text{GW}}$, with some differences in the noise-dominated regime, while the FCN shows an advantage in terms of the computation time. This is probably because the architecture of the FCN is simpler compared to that of the GNN, and the GNN requires handling graph structures, which can be more complex for simple mock datasets. This complexity makes the implementation and computation less efficient and less competitive.
@@ V Conclusions
Specifically, we analyze the performance of two NN architectures, an FCN and a GNN on a regression task to predict the value of the SGWB amplitude $\Omega_{\text{GW}}$. The results highlight distinct strengths and limitations for each architecture. The FCN consistently exhibits a plateau effect for small values across all tests, indicating its limitations in capturing the correct relationships in this range. This plateau persists even when the number of sources is varied or the datasets are masked. This behavior can be interpreted as a sign of reduced accuracy, particularly for small values of $\Omega_{\text{GW}}$. In contrast, the GNN tracks the ideal diagonal in the predictions with increasing dispersion for smaller $\Omega_{\text{GW}}$.
The GNN exhibits more stable behavior in terms of extracting the value of $\Omega_{\text{GW}}$, as the distribution of the predictions is closer to Gaussian, and the sensitivity scales with the theoretically predicted trend $\propto N_{s}$, with some exceptions for the large $N_{s}$ case. On the other hand, the FCN, while showing good sensitivity overall, saturates for specific configurations, but it demonstrates a significant advantage in terms of computation time. Our results suggest that the choice of architecture should be guided by the specific requirements of the task, balancing accuracy, stability, and computational efficiency.
NNs will become essential tools for analyzing the large volumes of upcoming astrometric data with accounting for systematic uncertainties. This paper provides the first step toward the new application of NNs to GW astrometry. In particular, the GNN implementation is a relatively unexplored field, and we expect that its advantages will become more pronounced when dealing with more complicated models of the dataset, particularly when applying the analysis to real data. While there are still challenges to be addressed, the promising results from this work highlight the potential of NNs in advancing GW astrometry.
@@ Acknowledgments
The authors would like to thank S. Ferraiuolo for discussions at an early stage of the project, and T. S. Yamamoto and H. Takahashi for their helpful discussions. They also acknowledge support from the research project PID2021-123012NB-C43 and the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033. The authors also acknowledge the use of the IFT Hydra cluster. M.C. acknowledges support from the “Ramón Areces” Foundation through the “Programa de Ayudas Fundación Ramón Areces para la realización de Tesis Doctorales en Ciencias de la Vida y de la Materia 2023”. G.M. acknowledges support from the Ministerio de Universidades through Grant No. FPU20/02857. S.J. acknowledges support from the Agence Nationale de la Recherche (ANR) under contract ANR-22-CE31-0001-01. S.K. is supported by the Spanish Atracción de Talento contract no. 2019-T1/TIC-13177 granted by Comunidad de Madrid, the I+D grant PID2020-118159GA-C42 funded by MCIN/AEI/10.13039/501100011033, the i-LINK 2021 grant LINKA20416 of CSIC, and Japan Society for the Promotion of Science (JSPS) KAKENHI Grant no. 20H01899, 20H05853, and 23H00110.

### v2

@@ IV Results
In Fig. 2, we present the test performance results for both FCN (left panel) and GNN (right panel). The networks can perform quite well for higher values of $\Omega_{\text{GW}}$, while the predictions start to become less precise for lower values when the noise starts to dominate. The vertical line is the theoretically estimated sensitivity for $\Omega_{\text{GW}}$ given by Eq. (3) included for reference. Note that this theoretical estimate assumes the same proper motion errors for all sources, whereas our mock data do not satisfy this assumption, as the proper motion errors depend on the magnitude of the sources. The results do not necessarily need to completely align with the theoretical line. As can be seen, by increasing the number of sources, the lines move toward the left, due to the inverse proportionality to the number of objects as seen in Eq. (3) and the NNs predictions also improve with an increasing number of sources.
@@
Note that the plot is shown in log scale, while our mock catalog dataset is generated with a linear-uniform distribution of $\Omega_{\text{GW}}$. We have tested both linear-uniform and log-uniform distribution of $\Omega_{\text{GW}}$ and found that NNs perform better with the linear-uniform distribution. Specifically, for the linear-uniform case, we obtain a reduced variance in the estimator (tighter scatter) and a more consistent linear trend across the tested range of $\Omega_{\text{GW}}$. It is worth noting that the choice of training distribution represents an implicit prior, as NNs effectively learn the probability density of the training data. Consequently, the network’s performance is intrinsically linked with these training settings, and the model may require retraining if the underlying physical prior for $\Omega_{\text{GW}}$ significantly deviates from our assumed distribution. Additionally, for the GNN, we evaluated the impact of varying the distance threshold values used to define the graph connectivity for a fixed number of sources. We observed that variations in the distance threshold did not lead to substantial differences in the trained performance of the GNN.
@@
In both plots, the predictions follow the diagonal (black dashed line), showing that both networks recover the expected scaling of the signal. However, below $10^{-1}$, the predictions are more scattered. These differences can be linked to the intrinsic design of each architecture. Note that both architectures are already optimized, and further changes in complexity do not lead to better results, but only to a higher computational cost. In particular, for the GNN this could indicate that the scatter is not only a function of the model depth but it might also come from intrinsic factors such as how the network aggregates and propagates information within the graph structure. Although deeper GNNs are better for capturing complex dependences, they also introduce challenges such as greater variability in predictions. These observations underscore the complex relationship between model design choices and performance outcomes, resulting in careful tuning of model architectures to balance precision and stability for specific tasks.
Figure 3: Test of trained FCN and GNN networks to new test datasets obtained by fixing $\Omega_{\text{GW}}$ to 0.03,0.3,1.00.03\,,0.3\,,1.0 values. Each violin illustrates the aggregated distribution of predictions across 100 independent realizations for each fixed value. Results are shown for different values of $N_{s}$, with corresponding MSE values reported in Table 1.
Finally, to estimate the predictive uncertainties and assess generalization, we apply the trained NNs to new unseen datasets and compute the MSE between the predictions and the true values. This allows us to test how reliably the networks perform beyond the training data, which is essential for evaluating their robustness in practical applications. The new mock datasets are made up of $100$ new independent realizations by fixing the values of $\Omega_{\text{GW}}$ to [0.03, 0.3, 1.0][0.03,\,0.3,\,1.0]. The violin plots in Fig. 3 visualize the distribution of the predictions across these 100 realizations for each fixed value. The corresponding MSE values are reported in Table 1.
MSE from FCN
$N_{s}$ $\Omega_{\text{GW}}=0.03$ $\Omega_{\text{GW}}=0.3$ $\Omega_{\text{GW}}=1.0$
$500$ $1.9626\times 10^{-2}$ $2.2690\times 10^{-2}$ $3.0275\times 10^{-2}$
$2000$ $3.3112\times 10^{-3}$ $3.6179\times 10^{-3}$ $9.9824\times 10^{-3}$
$12000$ $1.0028\times 10^{-3}$ $1.0326\times 10^{-3}$ $1.8717\times 10^{-3}$
MSE from GNN
$N_{s}$ $\Omega_{\text{GW}}=0.03$ $\Omega_{\text{GW}}=0.3$ $\Omega_{\text{GW}}=1.0$
$500$ $1.4777\times 10^{-1}$ $8.5298\times 10^{-2}$ $7.1031\times 10^{-2}$
$2000$ $8.3903\times 10^{-3}$ $8.2136\times 10^{-3}$ $2.9764\times 10^{-2}$
$12000$ $2.4736\times 10^{-3}$ $1.5479\times 10^{-3}$ $8.3370\times 10^{-3}$
Table 1: MSE values of predictions on the test dataset for both the FCN and the GNN. The comparison is shown across different prediction targets, specifically $\Omega_{\text{GW}}\in[0.03,0.3,1.0]$.
Comparison with MCMC
Figure 4: Comparison in logarithmic scale between MCMC (orange) and FCN (blue) predictions on $\Omega_{\text{GW}}$ values, uniformly distributed in the range $[0,1]$. The number of test samples is $1600$ for the case $N_{s}=2000$.
It is of particular interest to compare the sensitivity of our NNs with that of likelihood-based methods, since MCMC approaches represent the state of the art in the literature [23]. In Fig. 4, we show the comparison between the FCN and MCMC predictions across the entire original test dataset, which consists of $1600$ mock realization, for the representative case of $N_{s}=2000$. The MCMC algorithm used is the same as Ref. [23]. We restrict the comparison with MCMC to the FCN case.
The MCMC model predictions appear tightly clustered around the diagonal, indicating high accuracy with potentially lower uncertainty. Nevertheless, the FCN seems to be performing close to the traditional estimator, generally following the same trend, displaying a slightly wider spread. Moreover, though FCN predictions exhibit a somewhat broader spread, they still capture the main trend accurately and crucially in a fraction of the time, generating predictions in order of minutes, compared to approximately the order of days required by the MCMC method for the same task. This highlights a major practical advantage of ML that is significantly faster inference, making them especially valuable when rapid predictions are needed or when working with large datasets. However, unlike the MCMC approach, which naturally provides full posterior distributions and credible intervals, the FCN approach introduced here provides point estimates only.
@@
Figure 6: The variance of the residual distributions in Fig. 5 is plotted as a function of the number of sources. In the left panel, solid lines represent the results for the predictions in Fig. 5 for the FCN (yellow) and GNN (brown). Other curves correspond to the results obtained by varying the number of mock data. These predictions are generated using test datasets, which constitute $20\%$ of the original mock datasets. The original mock datasets consist of $5000$ (dotted lines), $8000$ (solid lines), and $16000$ (dash-dotted lines) samples, respectively. The blue dashed line represents the theoretical sensitivity estimate, where the statistical error decreases proportionally to $1/\sqrt{N_{s}}$, with $N_{s}$ being the number of sources. In the right panel, the variance of the residual distributions is shown by varying the range of $\Omega_{\text{GW}}$ sampling. The default range of $[0,1]$ is reduced to $[0,0.3]$ and further to $[0,0.1]$. The number of mock datasets is fixed at $8000$. Results are displayed only for the GNN.
@@
Figure 7: Predicted values of $\Omega_{\text{GW}}$ for the masked dataset are tested by changing the area of the mask. The number of sources is fixed to $N_{s}=2000$. The left panel represents the result of FCN, while the right panel the one from the GNN.
@@
Figure 8: The probability distributions of deviations between the true and predicted values for $\Omega_{\text{GW}}$ are plotted using the results obtained from the test dataset in Fig. 7, in the case of masked datasets. The cyan curve represents the Gaussian fit to the FCN results, while the blue curve represents the Gaussian fit to the GNN results.
Distributions of the variance and its dependence on the source number
To further analyze the performance of the models, we plot in Fig. 5 the distribution of the residuals, i.e., the difference between the predicted and true values. We observe that the variance of the FCN is generally smaller than that of the GNN. For reference, we also include the Gaussian fit of the histogram in the plot.
Furthermore, we observe a clear trend where the width of the distribution decreases as the number of sources increases. To investigate this trend more explicitly, Fig.6 shows the variance as a function of the number of sources $N_{s}$. For reference, we also plot a line proportional to $N_{s}^{-1/2}$.
The left panel compares the FCN and GNN results. Consistent with Fig. 5, the FCN exhibits a smaller variance compared to the GNN. To investigate this further, we conduct additional tests by varying the number of mock data $N$ plotted in the same figure based on the hypothesis that the requirement for more catalog realizations becomes increasingly critical as $N_{s}$ grows. We observe that the result is not affected when comparing the lines for $N=5000$ and $N=8000$.
In the right panel, we show that the range of $\Omega_{\text{GW}}$ for the mock data should be optimized based on the sensitivity.
Inhomogeneous case: masked samples
Above, we have demonstrated the application of NNs with simplified mock data settings. However, there are many complex factors to consider when handling real data, and this presents a much longer path to fully realizing the potential of NNs. We leave it for the future work, while here, to illustrate one example of the flexibility of NNs, we apply them to masked data and evaluate their performance.
We fix the number of sources to $N_{s}=2000$ and mask the regions of the dataset corresponding to the declination of $\delta=10^{\circ}$, $20^{\circ}$, $40^{\circ}$, and $60^{\circ}$. While we are aware that $\delta=60^{\circ}$ is a non-realistic value for the mask, it is used here solely for training performance purposes. The same NN architectures are trained on these masked subsets and the results are shown in Figs. 7 and 8.
Despite the variation in input data coverage, the overall behavior of the two architectures remains qualitatively consistent, as seen in Fig. 7. These results suggest that both networks remain robust under inhomogeneous sky coverage, with performance gradually degrading as masking increases. This behavior is evident in the distributions shown in Fig. 8, where the variance clearly rises with the mask size. The FCN tends to produce slightly narrower residual distributions, indicating marginally better performance in terms of accuracy, whereas the GNN displays a broader spread around the true values. However, one can notice from Fig. 8 that as the mask size increases, the FCN performance degrades more noticeably than that of the GNN. This suggests that the GNN, while less accurate in the absence of strong masking, is comparatively more robust to missing or more complicated datasets. Both networks successfully recover the dominant $\Omega_{\rm GW}$ signal, but the dispersion of predictions increases systematically as the masked region grows, in agreement with expectations.
Computation time
Given the expected number of sources available in future data, a reduction in computation time is primarily anticipated for the NNs. To address this, the computational time (including the training time) is plotted in Fig. 9 as a function of the number of sources. All these results were evaluated on a GPU cluster. [^1] For completeness, the specs of the GPU cluster are an INTEL ICE LAKE 6338N 2.20GHZ CPU, with 32 GB DDR4 RAM, and four NVIDIA A100 80 GB CoWoS HBM2 PCIe 4.0 GPUs, along with three HDDs of 4 TB each, and an Infiniband adapter for fast connectivity between the nodes. As observed, in both cases, the computation time increases logarithmically with the number of sources. The GNN exhibits higher computational time, which highlights the greater computational complexity inherent to the network architecture and the handling of more complex input data. In both cases, the behavior reflects a more accentuated computational cost for larger datasets.
Both networks are competitive in predicting the value of $\Omega_{\text{GW}}$, while the FCN shows an advantage in terms of the computation time. This is probably because of the architectures, and the GNN requires handling more complex graph structures. This complexity may make the implementation and computation less efficient and less competitive.
@@ V Conclusions
Specifically, we analyze the performance of two NN architectures, the FCN and GNN on a regression task to predict the value of the SGWB amplitude $\Omega_{\text{GW}}$. The results highlight distinct strengths and limitations for each architecture.
It is worth noting that the FCN is more reliable when a complete (original) dataset is available, as it exhibits a higher degree of precision in the prediction of the injected values, showing tighter clustering around the diagonal compared to those of the GNN, particularly as the number of samples increases. The corresponding histograms of the distribution of injection-predictions reinforce this, with the FCN distributions being narrower and centered closer to zero, indicating smaller prediction errors. This suggests that the FCN, when adequately trained in comprehensive data, can take advantage of the full feature space effectively to produce highly reliable results. However, when the mask is applied, the performance appears to degrade more noticeably. The scatter plots show a wider spread and a greater deviation from the diagonal, indicating a loss of predictive accuracy.
In contrast, the GNN demonstrates a certain robustness across all tested scenarios, including those with significant masking. Although the performance of the GNN might not be better than that of the FCN, it still consistently provides reasonable results even when substantial portions of the data are masked. This inherent robustness of GNNs is probably due to their ability to model relationships and dependences within the data. This structural understanding allows GNNs to make more informed predictions even when parts of the input are missing or obscured.
The observed differences in performance and robustness between the two networks are fundamentally rooted in their intrinsic architectural designs and how they process information. The FCN offers a significant advantage in terms of computation time. In general, our findings suggest that the choice of architecture should be guided by the specific requirements of the task, balancing accuracy, stability, and computational efficiency.
An interesting direction for future work could be to apply likelihood-free inference methods [51, 52], which are particularly well suited to cases where explicit likelihoods are intractable or the analysis is very complicated and computationally expensive. Traditional MCMC-based methods can be prohibitively slow despite their ability to naturally provide credible intervals. By adopting likelihood-free inference approaches based on posterior estimation, it becomes possible to generate full posterior distributions, thereby recovering the formal credible intervals as in standard Bayesian analyses, while keeping the computational efficiency of NNs. In the context of GW astronomy, such approaches have been employed to infer source properties of the compact binaries  [53, 54, 55, 56], and extended to the SGWB [57, 58], enabling a direct comparison with existing MCMC-based results in the literature.
NNs will become essential tools for analyzing the large volumes of upcoming astrometric data with accounting for systematic uncertainties. This paper provides the first step toward the new application of NNs to GW astrometry. While there are still challenges to be addressed, the promising results from this work highlight the potential of NNs in advancing GW astrometry.
@@ Acknowledgments
The authors would like to thank S. Ferraiuolo for discussions at an early stage of the project, and T. S. Yamamoto and H. Takahashi for their helpful discussions. They also acknowledge support from the Research Project No. PID2021-123012NB-C43 and the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No. CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033. The authors also acknowledge the use of the IFT Hydra cluster. M.C. acknowledges support from the Ramón Areces Foundation through the Programa de Ayudas Fundación Ramón Areces para la realización de Tesis Doctorales en Ciencias de la Vida y de la Materia 2023. G.M. acknowledges support from the Ministerio de Universidades through Grant No. FPU20/02857. S.J. acknowledges support from the Agence Nationale de la Recherche (ANR) under Contract No. ANR-22-CE31-0001-01. S.K. is supported by the Spanish Atracción de Talento Contract No. 2023-5A-TIC-28945 granted by Comunidad de Madrid, the I+D Grant No. PID2023-149018NB-C42 funded by MCIN/AEI/10.13039/501100011033, and the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant No. JP20H05853, No. JP23H00110, and No. JP24K00624.